Let S = {u1,u2,u3} be a basis for the vector space V. Show that T = {w1,w2,w3} is also



a basis for V, where w1 = u1 +u2 +u3, w2 = u2 +u3,w3 = u3.

1
Expert's answer
2022-04-07T14:33:54-0400

A vector space of dimension n has as many bases as there are linearly independent systems of n n -dimensional vectors. In other words, the basis of an n-dimensional space is any system of n independent vectors of this space.

We must show that number of vectors w1, w2, w3 is equal the dimension of V and these vectors are linearly independent.

The dimension of space V is 3 because there are 3 vectors in the basis S. T also contains three vectors, so it can be a basis for V if w1, w2, w3 are linearly independent.

We know that u1, u2, u3 are linearly independent because S is a basis for V. It means that a1u1 +a2u2 +a3u3=0 only when a1=0, a2=0 and a3=0. Vectors w1 = u1 +u2 +u3, w2 = u2 +u3, w3 = u3 are linear combination of u1, u2, u3.

Let w1, w2, w3 are linearly dependent. In this case linear combination b1w1+b2w2+b3w3=0 exists for nonzero b1, b2, b3.

If substitute w1 = u1 +u2 +u3, w2 = u2 +u3, w3 = u3:

b1(u1 +u2 +u3)+b2(u2 +u3)+b3(u3)=0

Open the brackets and groupe:

b1u1 +b1u2 +b1u3+b2u2 +b2u3+b3u3=0

b1u1 +(b1+b2)u2 +(b1+b2+b3)u3=0.

Last combination can not exist for nonzero b1, b2, b3 because u1, u2, u3 are linearly independent vectors.

Thus, w1, w2, w3 are linearly independent and T = {w1,w2,w3} is a basis for vector space V.